When we consider multiple random samples from a given population, the mean of these samples tend to be closer to the population mean

A> If sample size of these multiple samples are increased
B> If sample size of these multiple samples are decreased
C> Increasing or decreasing the sample size of these multiple samples makes no difference as the samples considered are true random samples
D> None of the above
1
Expert's answer
2020-03-25T11:47:25-0400

The correct answer is A) If sample size of these multiple samples are increased the mean of these samples tend to be closer to the population mean.

"E=z_{\alpha/2}\frac{\sigma}{\sqrt{n}}" is the margin of error.

We see that if sample size increases margin of error decreases.

From here it follows that "n=(\frac{z_{\alpha/2}\sigma}{E})^2."